# Electromagnetic Radiation from High-Energy Nuclear Collisions

### v1

@@ 2.2 Net photon yields
Figure 4: The yield distribution (top panels) of real photons shown with the appropriate experimental data measured in a $0-20\%$ centrality window. The contributions of the different processes discussed in the text are shown separately. (Left panel) Photons from Au + Au collisions at 200 $A$ GeV. The data are from the PHENIX PHENIX:2014nkk and STAR STAR:2016use Collaborations. (Right panel) Photons from Pb + Pb collisions at 2760 $A$ GeV. The data are from the ALICE Collaboration ALICE:2015xmh . The bottom panels show the ratio of data over the calculated result. Figure adapted from Gale:2021emg .
The sum of direct real photons (long dashes) is further decomposed into channels which include: photons from the first-impact collisions calculated with pQCD (“prompt”, short dashes); pre-hydrodynamics photons (“pre-equilibrium”, solid); thermal photons from the QGP and from the ensemble of composite hadrons (dot-dash). One observes that at RHIC, the pre-equilibrium, thermal, and pQCD photon contributions are comparable at $p_{T}\approx 3$ GeV, and at the LHC, for $p_{T}\approx 3.5$ GeV: the pre-equilibrium contribution represents $\gtrsim 30\%$ of the net photon yield and thus has a promising future as an observable signal. At the LHC energy shown in Fig. 4, the model and data agree within statistical uncertainties. Note in passing the unfortunate and enduring disparity between PHENIX and STAR results David:2019wpt .
Figure 5: The direct photon transverse momentum spectrum measured by the PHENIX, STAR, and ALICE Collaborations, divided by the calculations of Gale:2021emg . The figure is from Ref. Reygers:2022crp (Creative Common License, CC-BY 4.0).
@@ 2.3 Net photon flow
In the equation above, $v_{n}^{\gamma}$ is the photon flow coefficient, $v_{n}^{h}$ is that of charged hadrons, and $\Psi_{n}^{i}$ is the event plane angle of photons ($i=\gamma$) and hadrons ($i=h$). The hadrons are used to define a reference plane, and the photon momentum anisotropy is measured with respect to this plane Paquet:2015lta . Therefore, prior to the consideration of electromagnetic observables and of their momentum anisotropy, it is imperative that the hadronic sector of any calculation be in agreement with the corresponding data.
@@ 2.4 Dileptons
The evolution of thermal rates for dilepton emission Baier:1988xv ; Altherr:1989jc ; Aurenche:2002wq ; Laine:2013vma has proceeded in parallel with that of real photons, and the effect of viscosity corrections has been discussed and highlighted there as well Vujanovic:2016anq ; Vujanovic:2017psb ; Vujanovic:2019yih . As was the case for the emission of real photons, not all channels have been corrected for viscous effects. For example, low invariant mass ($M<m_{\phi}$) lepton pairs will receive a large contributions from baryons affecting $\Pi_{\mu\nu}$ in Eq. (1) Rapp:2009az ; estimates of viscous effects there are still being performed. In the QGP, the electromagnetic emissivity of an ensemble of hot partons is known today at NLO, and at finite baryon density, $\mu_{\rm B}\neq 0$ Churchill:2023vpt . Note that photon rates complete at ${\mathcal{O}}(\alpha_{s})$, $\mu_{\rm B}\neq 0$, and including the LPM effect had been established in Ref. Gervais:2012wd . Both of those will be treated in the future for non-equilibrium effects owing to viscosity.
@@
The dilepton-equivalent of the “cold photons” calculable in pQCD, are those originating from the Drell-Yan (DY) process Drell:1970wh . The calculations of DY cross sections have also been integrated in numerical packages such as Pythia Sjostrand:2006za , or the more specialized DYTurbo Camarda:2019zyx which performs a transverse-momentum resummation up to next-to-next-to-leading logarithmic accuracy, combined with fixed-order results at next-to-next-to-leading order (NNLO). Those two permit the evaluation of the irreducible background to a thermal dilepton signal.
@@
Even though the thermal dilepton yield stems from the same current-current correlator as for the real photons, measurements of dileptons offer in addition the potential to reveal the status of chiral symmetry restoration, as mentioned previously, as well as the hope to access an in-medium temperature that is impervious to kinematic distortions. The “dilepton thermometer” exists because the yield of lepton pairs can be calculated and measured as a function of their invariant mass $M^{2}~{}=~{}(p_{+}~{}+~{}p_{-})^{\mu}~{}(p_{+}~{}+~{}p_{-})_{\mu}$, whereas real photons are confined to the light cone and their spectrum depends on their momentum which is not a Lorentz invariant. For instance, photon spectra will experience Doppler shift: the local motion in the lab frame of a given cell will distort the fluid rest frame photon distribution. This makes photon and dilepton measurements complementary: in addition to the effective temperature information, the photon spectra will also inform the fluid dynamical modeling vanHees:2011vb ; Shen:2013vja ; Massen:2024pnj , while the dilepton spectra will offer a Doppler-free reading. Note however that measurements will still consist of lepton pairs emitted at all stages of Fig. 1, and over a distribution of cell temperatures.
@@
In terms of measurements, low and intermediate mass dileptons were previously analyzed by the pioneering DLS experiment DLS:1997kbk at the LBNL Bevalac, and at the CERN SPS by the CERES NA38/NA50, and Helios-3 Collaborations Specht:2007ez . In the low invariant mass region (LMR), $M\lesssim 1$ GeV, the dilepton spectrum measured with great precision by the NA60 Collaboration in collisions of In + In at $A$ 158 GeV reveals NA60:2008ctj a striking broadening of the $\rho$ meson spectral density, where coupling with baryon resonances plays a decisive role Rapp:1999ej . At RHIC, dileptons are investigated by the STAR Geurts:2012rv and PHENIX PHENIX:2005kfn Collaborations, and at the LHC by ALICE Feuillard:2022ket . For the collider conditions prevailing at RHIC, the STAR collaboration has reported dilepton measurements at various energies, namely BES-I and -II (Beam Energy Scan) and $\sqrt{s_{\rm NN}}$ = 200 GeV. The LMR measurements agree CJHP2024 with calculations based on the same spectral density integrated with a thermal fireball model previously found to match the one extracted by NA60 from the In + In data Huck:2014mfa , also supporting the scenario where $\rho$ meson broadening is driven by interaction with baryon resonances.
@@
The dilepton spectrum in the intermediate mass region (IMR), 1 GeV $<M\lesssim 3$ GeV has been the focus of some attention because of the possibility for dilepton measurements to reveal a possible restoration of chiral symmetry in a hot and dense strongly interacting system Rapp:1999ej , proceeding through a mixing of the vector and axial-vector spectral densities Geurts:2022xmk . This would however necessitate the observation of the pseudovector spectral density Kapusta:1993hq ; Hohler:2013eba ; Geurts:2022xmk , which has however remained elusive. Because a clear signal of this restoration is still ambiguous, the NA60+ Collaboration has put forward a new experiment proposal which features the search for chiral symmetry restoration effects through $\rho-a_{1}$ mixing, the study of the order of the phase transition at large baryochemical potential through the measurement of a caloric curve, and the onset of the deconfinement through the measurement of $J/\Psi$ suppression at the CERN SPS. The letter of intent NA60:2022sze states that chiral restoration would generate a 20-25% enhancement of the dilepton invariant mass spectrum over the case where the $\rho$ and the $a_{1}$ do not mix, around $M\sim$ 1 GeV. If approved, data taking could start in 2029.
@@
We leave the observation of chiral symmetry restoration as a tantalizing possibility for now and concentrate on temperature extractions, which has been the subject of previous phenomenological and experimental studies NA60:2008ctj ; Rapp:2014hha ; STAR:2024bpc . Recent analyses done using the multistage approach introduced earlier have also been completed Churchill:2023zkk ; Churchill:2023vpt . The first lesson from these recent papers is the importance of next-to-leading order (in the strong coupling) QCD contributions to the dilepton spectrum coming from the hot partonic medium, when compared to results obtained at leading order (LO). This is made clear in Fig. 7 which shows the thermal lepton pair production rate $\Gamma$ (@LO and @NLO) as a function of invariant mass for different medium temperatures. Clearly, NLO corrections have a considerable effect – qualitatively and quantitatively, depending on the temperature – at low ($M<1$ GeV) invariant mass, but also affect the intermediate mass (1 GeV $\lesssim M\lesssim 3$ GeV) region. Note that the need to go beyond the Born term for dilepton production at finite temperature had been known for some time Braaten:1989mz ; Wong:1991be . In the future, as was the case for some channels contributing to real photon emission, NLO dilepton emission rates will have to be corrected for non-equilibrium conditions; some of this has been done for LO rates Vujanovic:2016anq ; Vujanovic:2017psb ; Kasmaei:2018oag ; Vujanovic:2019yih .
@@
The emission of lepton pairs from the pre-hydro era had received some attention Martinez:2008di . Recent estimates based on using either kinetic theory or KøMPøST (as was done for photons) and including the contribution of the Drell-Yan leptons pairs have been completed Coquet:2021lca ; Garcia-Montero:2024lbl ; Wu:2024pba , and the pre-equilibrium dilepton spectra computed for LHC conditions in those two approaches have been found to have very similar characteristics JFHP2024 . The conclusion that follows from those studies is that a pre-equilibrium component can contribute measurably to the spectrum of dileptons in the intermediate mass region, at an LHC energy (5.02 TeV). Those dileptons will outshine those from the Drell-Yan process and the clarity of this signal will greatly benefit from the explicit measurement of lepton pairs coming from the simultaneous decay of open charm mesons, as planned by the next generation of experiments.
@@
Another aspect of electromagnetic signals from heavy-ion collisions that has yet to receive more attention is that of polarization. For heavy-ion collisions at lower energies that those studied at RHIC and at the LHC – those at SIS and at the CERN SPS – the angular lepton distributions reflecting the virtual photon polarization have been studied Speranza:2017hso ; Seck:2023oyt using coarse-grained transport models, and have been measured by the NA60 NA60:2008iqj and HADES Seck:2023oyt Collaborations. At those energies, the angular distribution of low invariant-mass dileptons is consistent with the presence of the resonances generated by the interactions of the composite hadrons, with little or no apparent sign of a QGP signal. In collider conditions, the polarization of the virtual photons produced in quark-antiquark annihilation at leading order has been studied in theory and used as a probe of the QGP momentum anisotropy Baym:2017qxy ; Coquet:2023wjk , and so has that of real photons Hauksson:2023dwh .
However, as was the case for the dilepton mass spectra, polarization studies of the QGP demand NLO corrections in the low and intermediate invariant mass regions: as $M\to 0$, the LO Born term $q\bar{q}\to\gamma^{*}$ has no phase space to produce a real photon. This is illustrated in the following: the virtual photon polarization can be observed via the angular distribution of final-state leptons in a given Lorentz frame. In the rest frame of the virtual photon, the helicity frame, the coefficient of the polar angle distribution of one of the lepton in the dilepton is $\lambda_{\theta}$ and can be shown to be related [^6] In a static medium, for lepton and antilepton with negligible rest mass. to the transverse (T) and longitudinal (L) components of the dilepton spectral density Wu:2024igf : $\lambda_{\theta}\sim\left(\rho_{\rm T}-\rho_{\rm L}\right)/\left(\rho_{\rm T}+\rho_{\rm L}\right)$. There again, one observes stark qualitative and quantitative differences between $\lambda_{\theta}^{\rm LO}$ and $\lambda_{\theta}^{\rm NLO}$. One can show that $\lim_{M\to 0}\lambda_{\theta}^{\rm NLO}=1$, because $\lim_{M\to 0}\rho_{\rm L}^{\rm NLO}=0$, as expected of a real photon. On the other hand, $\lim_{M\to 0}\lambda_{\theta}^{\rm LO}\neq 1$ Wu:2024igf . Finally, it is encouraging to note that, after properly averaging $\lambda_{\theta}$ over hydrodynamical fluid cells using a realistic space-time model, the characteristic profile survives the fluid evolution Wu:2024igf : clearly, electromagnetic polarization is another precious variable to signal the presence of a QGP.
@@ 3 Conclusions
It is evident that during the last five decades, the progress in theory and measurement of electromagnetic radiation emitted in relativistic heavy-ion collisions has been remarkable. This owes much to the advances in treating the electromagnetic emissivity in its own right, but this progress was realized concurrently with rapid developments in the modeling of relativistic nuclear collisions. It is exciting that a clear thermal signal has been seen in the data; its quantification requires accompanying theoretical modeling but the models have grown tremendously in maturity. Photons and dileptons can act as tomographic probes which can also inform our understanding of the space-time evolution of the interaction volume. This is indeed the era of multimessenger heavy-ion physics. What remains to be done in theory is integrating the many aspects we have seen in the paper into an end-to-end approach. This is not a small task but the roadmap is clear and if recent developments are any indication, its realization is within reach.
There are several aspects of the emission of electromagnetic in heavy-ion collisions that this article could not cover. A few examples are that of electromagnetic interferometry KumarSrivastava:1993mp ; Garcia-Montero:2019kjk , of real and virtual photon emission from small systems Shen:2016zpp , of radiation from from jet-medium interactions Srivastava:2002ic ; Yazdi:2022cuk , and the radiation of photons with very low momenta Bailhache:2024mck . Clearly, the study of electromagnetic emission from nuclear collisions is a rich and still evolving topic which is perhaps now entering its golden age.

### v2

@@ 2.2 Net photon yields
Figure 4: The yield distribution (top panels) of real photons shown with the appropriate experimental data measured in a $0-20\%$ centrality window. The contributions of the different processes discussed in the text are shown separately. (Left panel) Photons from Au + Au collisions at 200 $A$ GeV. The data are from the PHENIX PHENIX:2014nkk and STAR STAR:2016use Collaborations. (Right panel) Photons from Pb + Pb collisions at 2760 AA GeV. The data are from the ALICE Collaboration ALICE:2015xmh . The bottom panels show the ratio of data over the calculated result. Figure adapted from Gale:2021emg .
The sum of direct real photons (long dashes) is further decomposed into channels which include: photons from the first-impact collisions calculated with pQCD (“prompt”, short dashes); pre-hydrodynamics photons (“pre-equilibrium”, solid); thermal photons from the QGP and from the ensemble of composite hadrons (dot-dash). One observes that at RHIC, the pre-equilibrium, thermal, and pQCD photon contributions are comparable at $p_{T}\approx 3$ GeV, and at the LHC, for $p_{T}\approx 3.5$ GeV: the pre-equilibrium contribution represents $\gtrsim 30\%$ of the net photon yield and thus has a promising future as an observable signal. At the LHC energy shown in Fig. 4, the model and data agree within statistical uncertainties. At RHIC, there is unfortunately a long-standing difference between the photon spectra measured by STAR and PHENIX, which is so far not fully understood David:2019wpt .
Figure 5: The direct photon transverse momentum spectrum measured by the PHENIX, STAR, and ALICE Collaborations, divided by the calculations of Gale:2021emg . The figure is from Ref. Reygers:2022crp (Creative Common License, $C$-BY 4.0).
@@ 2.3 Net photon flow
In the equation above, $v_{n}^{\gamma}$ is the photon flow coefficient, $v_{n}^{h}$ is that of charged hadrons, and $\Psi_{n}^{i}$ is the event plane angle of photons ($i=\gamma$) and hadrons ($i=h$). The hadrons are used to define a reference plane, and the photon momentum anisotropy is measured with respect to this plane Paquet:2015lta . Therefore, prior to the consideration of electromagnetic observables and of their momentum anisotropy, it is verified that the theoretical modeling ot the hadronic sector is in agreement with the corresponding data.
@@ 2.4 Dileptons
The evolution of thermal rates for dilepton emission Baier:1988xv ; Altherr:1989jc ; Aurenche:2002wq ; Laine:2013vma has proceeded in parallel with that of real photons, and the effect of viscosity corrections has been discussed and highlighted there as well Vujanovic:2016anq ; Vujanovic:2017psb ; Vujanovic:2019yih . As was the case for the emission of real photons, not all channels have been corrected for viscous effects. For example, low invariant mass ($M<m_{\phi}$) lepton pairs will receive a large contributions from baryons affecting $\Pi_{\mu\nu}$ in Eq. (2) Rapp:2009az ; estimates of viscous effects there are still being performed. In the QGP, the electromagnetic emissivity of an ensemble of hot partons is known today at NLO, and at finite baryon density, $\mu_{\rm B}\neq 0$ Churchill:2023vpt . Note that QCD photon rates complete at ${\mathcal{O}}(\alpha_{s})$, $\mu_{\rm B}\neq 0$, and including the LPM effect had been established in Ref. Gervais:2012wd . Both of those will be treated in the future for non-equilibrium effects owing to viscosity.
@@
The dilepton-equivalent of the “cold photons” calculable in pQCD, are those originating from the Drell-Yan (DY) process Drell:1970wh . The calculations of DY cross sections have also been integrated in numerical packages such as Pythia Sjostrand:2006za , or the more specialized DYTurbo Camarda:2019zyx which performs a transverse-momentum resummation up to next-to-next-to-leading logarithmic accuracy, combined with fixed-order results at next-to-next-to-leading order (NNLO). Those two permit the evaluation of the irreducible background to the thermal dilepton signal.
@@
Even though the thermal dilepton yield stems from the same current-current correlator as for the real photons, measurements of dileptons offer in addition the potential to reveal the effects of chiral symmetry restauration Rapp:1999ej as well as the hope to access an in-medium temperature that is impervious to kinematic distortions. The “dilepton thermometer” exists because the yield of lepton pairs can be calculated and measured as a function of their invariant mass $M^{2}~=~(p_{+}~+~p_{-})^{\mu}~(p_{+}~+~p_{-})_{\mu}$, whereas real photons are confined to the light cone and their spectrum depends on their momentum which is not a Lorentz invariant. For instance, photon spectra will experience Doppler shift: the local motion in the lab frame of a given cell will distort the fluid rest frame photon distribution. This makes photon and dilepton measurements complementary: in addition to the effective temperature information, the photon spectra will also inform the dynamical modeling vanHees:2011vb ; Shen:2013vja ; Massen:2024pnj , while the dilepton spectra will offer a Doppler-free reading. Note however that measurements will still consist of lepton pairs emitted at all stages of Fig. 1, and over a distribution of cell temperatures.
@@
In terms of measurements, low and intermediate mass dileptons were previously analyzed by the pioneering DLS experiment DLS:1997kbk at the LBNL Bevalac, and at the CERN SPS by the CERES, NA38/NA50, and Helios-3 Collaborations Specht:2007ez . In the low invariant mass region (LMR), $M\lesssim 1$ GeV, the dilepton spectrum measured with great precision by the NA60 Collaboration in collisions of In + In at $A$ 158 GeV reveals NA60:2008ctj a striking broadening of the $\rho$ meson spectral density, where coupling with baryon resonances plays a decisive role Rapp:1999ej . At RHIC, dileptons are investigated by the STAR Geurts:2012rv and PHENIX PHENIX:2005kfn Collaborations, and at the LHC by ALICE Feuillard:2022ket . For the collider conditions prevailing at RHIC, the STAR collaboration has reported dilepton measurements at various energies, namely BES-I and -II (Beam Energy Scan) and $\sqrt{s_{\rm NN}}$ = 200 GeV. The LMR measurements agree CJHP2024 with calculations based on the same spectral density integrated with a thermal fireball model previously found to match the one extracted by NA60 from the In + In data Huck:2014mfa , also supporting the scenario where $\rho$ meson broadening is driven by interaction with baryon resonances.
@@
The dilepton spectrum in the intermediate mass region (IMR), 1 GeV $M\lesssim 3$ GeV has been the focus of some attention because of the possibility for dilepton measurements to reveal a possible restauration of chiral symmetry in a hot and dense strongly interacting system Rapp:1999ej , proceeding through a mixing of the vector and axial-vector spectral densities Geurts:2022xmk . This would necessitate the observation of the pseudovector spectral density Kapusta:1993hq ; Hohler:2013eba ; Geurts:2022xmk , which has however remained elusive. Because a clear signal is still ambiguous, the NA60+ Collaboration has put forward a new experiment proposal which features the search for chiral symmetry restauration effects, the study of the order of the phase transition at large baryochemical potential through the measurement of a caloric curve, and the onset of the deconfinement through the measurement of $J/\psi$ suppression at the CERN SPS. The letter of intent NA60:2022sze states that chiral restauration through $\rho-a_{1}$ mixing would generate a 20-25% enhancement of the dilepton invariant mass spectrum over the case where the $\rho$ and the $a_{1}$ do not mix in the dilepton invariant mass region around $M\sim$ 1 GeV. If approved, data taking could start in 2029.
@@
We leave the observation of chiral symmetry restauration as a tantalizing possibility for now and concentrate on temperature extractions, which has been the subject of previous phenomenological and experimental studies NA60:2008ctj ; Rapp:2014hha ; STAR:2024bpc . Recent analyses done using the multistage approach introduced earlier have also been completed Churchill:2023zkk ; Churchill:2023vpt . The first lesson from these recent papers is the importance of next-to-leading order (in the strong coupling) QCD contributions to the dilepton spectrum coming from the hot partonic medium, when compared to results obtained at leading order (LO). This is made clear in Fig. 7 which shows the thermal lepton pair production rate $\Gamma$ (@LO and @NLO) as a function of invariant mass for different medium temperatures. Clearly, NLO corrections have a considerable effect – qualitatively and quantitatively, depending on the temperature – at low ($M<1$ GeV) invariant mass, but also affect the intermediate mass (1 GeV $\lesssim M\lesssim 3$ GeV) region. Note that the need to go beyond the Born term for dilepton production at finite temperature had been known for some time Braaten:1989mz ; Wong:1991be . In the future, as was the case for some channels contributing to real photon emission, NLO dilepton emission rates will have to be corrected for non-equilibrium conditions; some of this has been done for LO rates Vujanovic:2016anq ; Vujanovic:2017psb ; Kasmaei:2018oag ; Vujanovic:2019yih .
@@
The emission of lepton pairs from the pre-hydro era has also received some attention Martinez:2008di . Recent estimates based on using either kinetic theory or KøMPøST (as was done for photons) and including the contribution of the Drell-Yan leptons pairs have been completed Coquet:2021lca ; Garcia-Montero:2024lbl ; Wu:2024pba , and the pre-equilibrium dilepton spectra computed for LHC conditions in those two approaches have been found to have very similar characteristics JFHP2024 . The conclusion that follows from those studies is that a pre-equilibrium component can contribute measurably to the spectrum of dileptons in the intermediate mass region at LHC energy (5.02 TeV). Those dileptons will outshine those from the Drell-Yan process; however the unambiguous interpretation of this signal will greatly benefit from the explicit measurement of lepton pairs coming from the simultaneous decay of open charm mesons, as planned by the next generation of experiments.
@@
Another aspect of electromagnetic signals from heavy-ion collisions that is bound to receive more attention is that of polarization. For heavy-ion collisions at lower energies that those studied at RHIC and at the LHC – those at SIS and at the CERN SPS – the angular lepton distributions reflecting the virtual photon polarization have been studied Speranza:2017hso ; Seck:2023oyt using coarse-grained transport models, and have been measured by the NA60 NA60:2008iqj and HADES Seck:2023oyt Collaborations. At those energies, the angular distribution of low invariant-mass dileptons is consistent with the presence of the resonances generated by the interactions of the composite hadrons, with little or no apparent sign of a QGP signal. In collider conditions, the polarization of the virtual photons produced in quark-antiquark annihilation at leading order has been studied in theory and used as a probe of the QGP momentum anisotropy Baym:2017qxy ; Coquet:2023wjk , and so has that of real photons Hauksson:2023dwh .
However, as was the case for the dilepton mass spectra, polarization studies of the QGP demand NLO corrections in the low and intermediate invariant mass regions: as $M\to 0$, the LO Born term $q\bar{q}\to\gamma^{*}$ has no phase space to produce a real photon. This is illustrated in the following: the virtual photon polarization can be observed via the angular distribution of final-state leptons in a given Lorentz frame. In the rest frame of the virtual photon, the helicity frame, the coefficient of the polar angle distribution of one of the leptons in the dilepton is $\lambda_{\theta}$ and can be shown to be related [^6] In a static medium, for lepton and antilepton with negligible rest mass. to the transverse (T) and longitudinal (L) components of the dilepton spectral density Wu:2024vyc : $\lambda_{\theta}\sim\left(\rho_{\rm T}-\rho_{\rm L}\right)/\left(\rho_{\rm T}+\rho_{\rm L}\right)$. There again, one observes stark qualitative and quantitative differences between $\lambda_{\theta}^{\rm LO}$ and $\lambda_{\theta}^{\rm NLO}$. One can show that $\lim_{M\to 0}\lambda_{\theta}^{\rm NLO}=1$, because $\lim_{M\to 0}\rho_{\rm L}^{\rm NLO}=0$, as expected of a real photon. On the other hand, $\lim_{M\to 0}\lambda_{\theta}^{\rm LO}\neq 1$ Wu:2024vyc . Finally, it is encouraging to note that, after properly averaging $\lambda_{\theta}$ over hydrodynamical fluid cells using a realistic space-time model, the characteristic profile survives the fluid evolution Wu:2024vyc : clearly, electromagnetic polarization is another precious variable to signal the presence of, and to characterize, the QGP.
@@ 3 Conclusions
It is evident that during the last five decades, the progress in theory and measurement of electromagnetic radiation emitted in relativistic heavy-ion collisions has been striking. This owes much to advances in the treatment of electromagnetic emissivity, but this progress was realized concurrently with rapid developments in the modeling of relativistic nuclear collisions. It is exciting that data has suggested the existence of a clear thermal signal NA60:2007lzy which is on the verge of precise quantification Churchill:2023vpt . Photons and dileptons can act as tomographic probes which can also inform our understanding of the space-time evolution of the interaction volume. Given the many sources of electromagnetic quanta and the sophistication of modern multistage modeling of relativistic heavy-ion collisions, the fact that photon spectrum calculations and measurements show some level of agreement – even if improvement is warranted – is nothing short of remarkable. This is indeed the era of multimessenger heavy-ion physics. What remains to be done in theory is integrating the many stages we have described in this paper into an end-to-end approach. This is not a small task but the roadmap is clear and if recent developments are any indication, its realization is within reach. While experimental measurements of electromagnetic radiation are complicated because of small signal strength and large backgrounds, the community has spent significant effort to systematically include them alongside hadronic observables.
There are several aspects of the emission of electromagnetic radiation in heavy-ion collisions that this article could not cover. A few examples are that of electromagnetic interferometry KumarSrivastava:1993mp ; Garcia-Montero:2019kjk , of real and virtual photon emission from small systems Shen:2016zpp , of radiation from from jet-medium interactions Srivastava:2002ic ; Yazdi:2022cuk , and the radiation of photons with very low momenta Bailhache:2024mck . Clearly, the study of electromagnetic emission from nuclear collisions is a rich and still evolving topic which is perhaps only now entering its golden age.